Fultz Company has accumulated the following budget data for the year 2020.
1. Sales: 31,380 units, unit selling price $90.
2. Cost of one unit of finished goods: direct materials 1 pound at $5 per pound, direct labor 3 hours at $12 per hour, and manufacturing overhead $6 per direct labor hour.
3. Inventories (raw materials only): beginning, 10,370 pounds; ending, 15,270 pounds.
4. Selling and administrative expenses: $170,000; interest expense: $30,000.
5. Income taxes: 30% of income before income taxes.

Expand question part

(a)

Collapse question part

(b)

Prepare a budgeted multiple-step income statement for 2020.
FULTZ COMPANY
Budgeted Income Statement

For the Year Ending December 31, 2020For the Quarter Ending December 31, 2020December 31, 2020

Total Operating ExpensesSelling and Administrative ExpensesPurchasesSalesBeginning InventoryCost of Goods SoldOperating ExpensesEnding InventoryIncome Tax ExpenseGross ProfitNet Income / (Loss)Income before Income TaxesIncome from Operations

$

Ending InventorySalesPurchasesSelling and Administrative ExpensesTotal Operating ExpensesBeginning InventoryCost of Goods SoldGross ProfitIncome before Income TaxesIncome from OperationsIncome Tax ExpenseNet Income / (Loss)Operating Expenses

Beginning InventoryCost of Goods SoldTotal Operating ExpensesGross ProfitSelling and Administrative ExpensesIncome Tax ExpenseIncome before Income TaxesPurchasesIncome from OperationsSalesNet Income / (Loss)Ending InventoryOperating Expenses

Income before Income TaxesOperating ExpensesNet Income / (Loss)PurchasesSalesTotal Operating ExpensesSelling and Administrative ExpensesBeginning InventoryCost of Goods SoldIncome from OperationsEnding InventoryGross ProfitIncome Tax Expense

Net Income / (Loss)Cost of Goods SoldSelling and Administrative ExpensesOperating ExpensesSalesTotal Operating ExpensesBeginning InventoryPurchasesEnding InventoryGross ProfitIncome before Income TaxesIncome from OperationsIncome Tax Expense

Beginning InventoryIncome from OperationsInterest ExpensePurchasesTotal Operating ExpensesCost of Goods SoldEnding InventorySalesSelling and Administrative ExpensesGross ProfitIncome before Income TaxesIncome Tax ExpenseNet Income / (Loss)

Income before Income TaxesCost of Goods SoldIncome from OperationsIncome Tax ExpenseNet Income / (Loss)SalesSelling and Administrative ExpensesTotal Operating ExpensesOperating ExpensesPurchasesGross ProfitEnding InventoryBeginning Inventory

Ending InventoryIncome from OperationsIncome Tax ExpenseGross ProfitNet Income / (Loss)Total Operating ExpensesOperating ExpensesBeginning InventoryPurchasesSalesSelling and Administrative ExpensesIncome before Income TaxesCost of Goods Sold

Cost of Goods SoldEnding InventoryGross ProfitIncome before Income TaxesTotal Operating ExpensesOperating ExpensesBeginning InventoryIncome from OperationsIncome Tax ExpenseSelling and Administrative ExpensesNet Income / (Loss)PurchasesSales

$

Click if you would like to Show Work for this question:

Open Show Work
Answer #1
FULTZ COMPANY
Budgeted Income Statement
For the Year Ending December 31, 2020
Sales 2824200 =31380*90
Cost of goods sold 1851420 =31380*59
Gross Profit 972780
Selling and Administrative Expenses 170000
Income from Operations 802780
Interest Expense 30000
Income before Income Taxes 772780
Income Tax Expense 231834
Net Income / (Loss) 540946
Workings:
Cost of goods sold per unit:
Direct materials 5 =1*5
Direct labor 36 =3*12
Manufacturing overhead 18 =3*6
Cost of goods sold per unit 59
Income Tax Expense 231834 =772780*30%